# Revela: Dense Retriever Learning via Language Modeling

### v1

@@ 7 Conclusion
Appendix A Experiments
A.1 Evaluation Benchmarks
@@
A.2 Examples of the Training Corpus
Fig. 8 presents an example batch containing 16 chunks from two topics. When training, the chunks will be randomly shuffled.
@@ A.2 Examples of the Training Corpus
A.3 REPLUG
Self-supervised batch loss.
@@ A.3 REPLUG
Experimental setups
Both of temperatures, $\tau_{\!lm}$ and $\tau_{r}$, are set as 0.001. REPLUG is training on the same datasets with Revela, with the learning rate $5e^{-4}$ and the training steps 4500. All other experimental setups are identical to Revela.
A.4 Checkpoints
@@
LLaMA-3.2-1B meta-llama/Llama-3.2-1B Llama 3.x Community License
@@
A.5 Implementation of Reciprocal Rank Fusion
To combine the strengths of BM25 and Revela, we employ Reciprocal Rank Fusion (RRF) [42]. Given their ranked lists $R_{1}$ and $R_{2}$ for the same query, respectively, RRF assigns a fused score to every document $d$ according to
$\text{score}(d)\;=\;\frac{1}{k+\mathrm{rank}_{R_{1}}(d)}+\frac{1}{k+\mathrm{rank}_{R_{2}}(d)},$
where $\mathrm{rank}_{R_{i}}(d)$ is the $1$-based position of $d$ in run $R_{i}$ (undefined documents are ignored) and $k$ is a constant that controls how quickly the contribution of lower-ranked documents decays; we follow common practice and set $k=60$. After summing the two reciprocal ranks, the documents are re-sorted by this fused score, and the top 200 are retained as the final ranking delivered to RRF.
Appendix B Analysis
B.1 Analysis on the Size of LMs
Tables 5 and 6 present an ablation study analyzing the impact of differently sized LMs.
@@ B.1 Analysis on the Size of LMs
Table 5: NDCG@10 (%) on BEIR for three model sizes (Revela-135M, Revela-500M, Revela-1B), with LM sizes in ascending order. Each triplet of columns highlights the highest value in bold.
Revela-135M Revela-500M Revela-1B
@@
ArguAna 36.5 39.6 39.4 38.0 41.5 41.4 41.0 42.9 45.1
C-FEVER 14.4 16.6 15.6 17.8 18.4 13.5 14.4 15.1 15.6
DBPedia 17.5 17.0 18.3 24.2 22.7 21.4 24.9 26.4 27.6
FEVER 26.3 32.4 35.1 52.9 55.6 50.9 58.2 56.2 60.4
FiQA2018 16.9 19.2 19.1 26.4 24.8 26.6 29.2 27.9 29.6
HotpotQA 41.7 41.1 38.5 54.7 53.9 50.6 58.4 58.7 56.0
NFCorpus 20.0 22.3 23.5 27.3 27.3 26.5 26.3 28.9 27.2
NQ 22.4 21.0 21.2 34.5 32.2 29.8 36.8 34.9 33.9
QuoraRetrieval 79.2 79.8 81.0 83.4 83.5 83.1 82.4 83.3 83.5
SCIDOCS 10.8 11.3 12.1 15.3 15.1 14.8 15.9 16.0 16.4
SciFact 54.1 57.9 57.9 65.5 65.4 65.3 71.6 71.8 71.8
TRECCOVID 58.5 60.9 60.2 68.1 68.8 58.4 66.2 65.4 59.8
Touche2020 11.6 14.9 14.4 27.4 21.5 21.1 26.2 25.0 23.3
Mean 31.5 33.4 33.6 41.2 40.8 38.7 42.4 42.5 42.3
@@
Revela-135M Revela-500M Revela-1B

### v2

@@ Acknowledgments
Appendix A Illustration
To illustrate the training dynamics, we visualize both the in-batch attention computed by the retriever and the LM loss, for example, in Fig. 6. We use LLaMA-3.2‑1B for both the retriever and the LM, and compare at the initial checkpoint, 100 steps, and 200 steps. When the retrievers can model the semantics, the NTP loss decreases as shown in Fig. 6. In this example, (Blue, Yellow) and (Red, Purple) are semantically relevant pairs, where the former is related to biographical information of Aaron, while the latter is related to Aaron’s priesthood and religious duties
Figure 6: Example of training dynamics of Revela. The related patterns in red and purple sequences are highlighted in bold, underline, and italic.
Appendix B Experiments
B.1 Evaluation Benchmarks
@@
BRIGHT is a benchmark for reasoning-intensive retrieval across diverse domains, including StackExchange forums, coding tasks, and theorem-based math. It emphasizes cases with little lexical overlap between queries and relevant documents, exposing the limitations of existing models. Table 4 presents the specific subtaks included in BRIGHT with the corresponding descriptions.
Table 4: BRIGHT Benchmark Tasks. Abbreviations and descriptions.
Abbrev. Description
Bio. StackExchange Biology questions with external web evidence
Earth. StackExchange Earth Science questions with linked resources
Econ. StackExchange Economics questions requiring theoretical support
Psy. StackExchange Psychology questions with cited references
Rob. StackExchange Robotics questions involving technical docs
Stack. StackOverflow programming questions with linked web pages
Sus. StackExchange Sustainable Living questions with supporting docs
Leet. LeetCode problems; retrieve similar problems/solutions
Pony. Pony language coding tasks with syntax documentation
AoPS Math Olympiad problems; retrieve others using same skill
TheoQ. TheoremQA: retrieve problems applying the same theorem
TheoT. TheoremQA: retrieve theorems from ProofWiki relevant to query
B.2 Examples of the Training Corpus
Fig. 7 presents an example batch containing 16 chunks from two topics. The sentences are chunked by NLTK, [^5] https://www.nltk.org/ while the maximum length of a chunk is limited to 120 words. When training, the chunks will be randomly shuffled within the batch.
@@ B.2 Examples of the Training Corpus
B.3 REPLUG
Self‑supervised batch loss.
@@ B.3 REPLUG
In our experiments, both of temperatures, $\tau_{\!lm}$ and $\tau_{r}$, are set as 0.001. REPLUG is training on the same datasets with Revela, with the learning rate $5e^{-4}$ and the training steps 4500. All other experimental setups are identical to Revela.
B.4 Checkpoints
@@
UniXcoder (UniX) microsoft/unixcoder-base Apache-2.0
@@
E5-large-unsupervised (E5-PTlarge) intfloat/e5-large-unsupervised MIT
BGE-m3 BAAI/bge-m3 MIT
E5-Mistral-7B-Instruct intfloat/e5-mistral-7b-instruct MIT
@@
LLaMA-3.2-1B meta-llama/Llama-3.2-1B Llama 3.2 Community License Agreement
LLaMA-3.2-3B meta-llama/Llama-3.2-3B Llama 3.2 Community License Agreement
LLaMA-3.1-8B meta-llama/Llama-3.1-8B Llama 3.1 Community License Agreement
@@
B.5 Experimental setups
Model Architecture
When trained on Wikipedia, Revela prevents any single token from dominating the attention by scaling the output $\mathrm{b}^{l}_{ij}$ with the norm of $V_{j}^{e}$, thereby encouraging the cross-document attention to focus on sequence-level semantics (Izacard et al., 2023). This operation, referred to as V-normalization, is computed as
where $\epsilon$ is a small constant for numerical stability, set as $1\mathrm{e}{-6}$ in our experiment. When trained on code-related corpus and evaluated on CoIR, Revela performs better without V-normalization. We will leave the exploration of more variants of the architecture design in the future.
Similarity Calculation within Chunks
When calculating the similarity between chunks derived from Wikipedia, we follow REPLUG, using only the first half of the chunk to compute similarity with other chunks. This design is motivated by the inferential semantics inherent in natural language retrieval tasks. In contrast, for code retrieval tasks, we retain the full chunk when computing similarity, as these tasks rely more heavily on precise semantic matching compared to natural language.
B.6 Supplementary Results
Tab. 6 and Tab. 7 present the per-task results in BEIR and BRIGHT, including Revela and other baselines.
Table 6: Performance of unsupervised/self-supervised retriever models on BEIR datasets (NDCG@10, %). Bold marks the best score per dataset among unsupervised methods.
Dataset BM25 Contriever LaPraDor Revela E5 REPLUG Revela REPLUG Revela REPLUG Revela Revela-code
Model Size 0.1B 0.1B 0.1B 0.3B 0.5B 0.5B 1B 1B 3B 3B 3B
ArguAna 48.7 44.3 44.6 39.0 44.4 36.4 41.1 39.4 44.6 38.0 45.3 44.3
ClimateFEVER 13.6 7.2 12.2 15.3 15.7 16.2 13.8 15.0 15.8 13.0 16.6 18.6
DBPedia 29.9 27.0 25.0 18.3 37.1 18.0 21.3 19.5 27.6 23.4 27.1 30.9
FEVER 48.1 27.2 33.6 33.7 68.6 50.4 51.1 51.6 61.7 54.3 62.9 66.3
FiQA2018 25.1 12.4 19.8 19.2 43.2 19.9 27.2 22.7 30.4 24.8 34.3 33.2
HotpotQA 56.9 41.0 30.4 38.5 52.2 35.6 50.6 42.4 56.0 33.9 58.8 57.8
NFCorpus 32.1 27.1 30.4 23.6 33.7 26.9 26.8 26.7 27.2 26.9 33.0 32.9
NQ 28.5 18.1 18.0 21.2 41.7 26.7 29.8 27.8 33.9 38.1 40.8 41.0
QuoraRetrieval 80.4 83.4 78.7 81.0 81.0 78.4 83.2 82.4 83.5 83.3 83.8 85.6
SCIDOCS 15.8 10.9 13.4 12.1 21.8 13.6 14.8 14.5 16.3 15.3 17.6 18.8
SciFact 68.7 59.1 49.9 57.9 72.3 65.0 66.0 67.6 71.9 73.7 73.3 72.0
TRECCOVID 62.2 18.2 22.9 60.6 61.8 44.9 58.4 39.4 60.1 46.2 66.7 60.7
Touche2020 33.1 7.2 8.9 12.4 19.8 11.8 19.9 14.7 25.7 1.5 26.9 30.5
Mean 41.8 29.5 29.8 33.3 45.6 34.2 38.8 35.7 42.7 36.3 45.2 45.6
Table 7: Performance on BRIGHT (nDCG@10, %). Bold marks the best performance.
Dataset BM25 E5-PT Revela Revela REPLUG Revela E5-Mistral Revela Cohere OpenAI Voyage
Model Size 0.3B 0.5B 1B 3B 3B 7B 8B
Biology 18.9 18.9 17.0 17.1 7.1 25.3 18.6 24.4 18.7 23.3 23.1
Earth Science 27.2 29.2 28.8 29.1 13.1 40.0 26.0 40.9 28.4 26.7 25.4
Economics 14.9 10.3 13.9 13.2 12.9 17.1 15.5 21.7 20.4 19.5 19.9
Psychology 12.5 17.6 13.2 14.9 15.1 21.4 15.8 24.9 21.6 27.6 24.9
Robotics 13.6 7.0 9.1 11.0 7.6 15.5 16.3 14.6 16.3 12.8 10.8
StackOverflow 18.4 7.4 9.0 13.2 6.0 14.0 11.2 15.0 18.3 14.3 16.8
Sustainable 15.0 14.1 13.9 15.4 11.6 19.1 18.1 19.7 17.6 20.5 15.4
Pony 7.9 1.8 3.9 3.0 1.1 6.5 4.9 2.0 1.9 2.4 1.5
LeetCode 24.4 13.9 16.9 16.0 15.3 16.0 28.7 18.4 26.8 23.6 30.6
AoPS 6.2 1.4 4.6 2.7 5.6 3.8 7.1 3.5 6.3 8.5 7.5
ThmQA-Thm 4.9 4.9 5.6 7.3 8.5 9.7 26.8 11.8 7.2 11.7 11.6
ThmQA-Q 10.4 10.1 14.0 14.6 14.5 14.6 26.1 16.1 15.7 23.5 27.4
Mean 14.5 11.4 12.5 13.1 9.9 16.9 17.9 17.8 16.6 17.9 17.9
Appendix C Analysis
C.1 Experimental Setups of Contriever Training
Our Contriever implementation fine-tunes a LoRA-adapted LLaMA-3.2-1B retriever on a 500k Wikipedia and 359k code dataset, using contrastive learning with a temperature of 0.01, EOS pooling, and query/passage prefixes as ”Query: ” and ”Passage: ”, respectively. The LoRA rank is 256, consistent with Revela. The training is optimized with DeepSpeed ZeRO-3, bfloat16 precision, gradient checkpointing, and an effective batch size of 256. The number of negatives for each query is 15, which is consistent with the popular contrastive learning setup such as MsMarco. [^6] https://huggingface.co/datasets/Tevatron/msmarco-passage-aug
The per-task results in Tab. 2 on CoIR and BEIR are displayed in Tab. 8 and Tab. 9. For in-domain evaluation, Revela outperforms the contrastive-learning counterpart on both BEIR and CoIR. Though the performance on BEIR is quite close, Revela demonstrate better generalization: it outperforms Contriever on tasks like FiQA, SciFact, and TRECCOVID, while Contriever exhibits better performance on ClimateFEVER or FEVER, which use Wikipedia as the corpus. For out-of-domain tasks, Revela significantly surpasses Contiever, highlighting the advantage over traditional contrastive learning paradigm.
Table 8: Performance on CoIR (NDCG@10, %). Bold marks the best score among the models.
Dataset Revela-wiki1B Contriever-wiki1B Revela-code1B Contriever-code1B
Apps 9.1 11.3 19.4 24.4
CosQA 22.6 24.3 30.2 32.4
ST2SQL 55.8 46.3 55.0 44.8
SN 54.2 49.8 64.0 41.5
SNCC 69.2 57.4 70.0 52.3
TransC 74.4 76.3 81.1 79.6
TransDL 34.7 33.8 34.2 30.8
SOQA 70.9 81.2 85.7 91.0
F-ST 70.6 65.5 76.2 66.0
F-MT 70.7 56.6 70.4 58.5
Mean 53.2 50.3 58.6 52.1
Table 9: Performance of unsupervised/self-supervised retriever models on BEIR datasets (NDCG@10, %). Bold marks the best score per dataset.
Dataset Revela-wiki1B Contriever-wiki1B Revela-code1B Contriever-code1B
ArguAna 44.6 48.9 46.9 42.0
ClimateFEVER 15.8 29.0 14.9 13.7
DBPedia 27.6 31.6 21.8 14.6
FEVER 61.7 68.6 51.4 41.2
FiQA2018 30.4 27.7 33.3 31.7
HotpotQA 56.0 47.1 46.4 17.6
NFCorpus 27.2 34.0 27.3 30.5
NQ 33.9 28.9 29.5 15.0
QuoraRetrieval 83.5 77.5 84.3 76.9
SCIDOCS 16.3 21.3 16.3 19.9
SciFact 71.9 63.9 66.4 61.5
TRECCOVID 60.1 55.5 54.2 39.5
Touche2020 25.7 17.4 21.8 15.2
Mean 42.7 42.4 39.6 32.3
C.2 Analysis on Batch sizes
Tab. 10 and Tab. 11 present the performance of Revela with different batch sizes on CoIR and BEIR, respectively. Generally, with a larger batch size, Revela’s performance will be increased.
Table 10: Retrieval performance (%, NDCG@10) across the 10 CoIR tasks. We report three encoder sizes (135M, 500M, 1B) and three batch sizes (bs4, bs8, bs16).
Revela0.1B Revela0.5B Revela1B
Dataset bs4 bs8 bs16 bs4 bs8 bs16 bs4 bs8 bs16
AppsRetrieval 4.8 6.4 8.2 11.8 16.5 20.5 17.1 16.2 19.4
CosQA 25.0 25.4 26.2 27.5 28.4 27.5 28.0 28.1 30.2
SyntheticText2SQL 42.8 44.0 45.7 50.9 52.5 53.7 53.4 51.4 55.0
COIRCodeSearchNetRetrieval 39.5 46.5 49.9 57.0 52.9 57.9 59.3 57.9 64.0
CodeSearchNetCCRetrieval 56.0 60.4 63.4 62.5 63.7 68.0 69.0 69.2 70.0
CodeTransOceanContest 63.4 68.1 70.9 71.6 74.0 77.6 75.6 77.7 81.1
CodeTransOceanDL 34.6 34.9 34.6 33.8 34.0 35.4 34.4 34.3 34.2
StackOverflowQA 57.1 62.8 69.2 73.8 76.6 82.5 77.6 82.5 85.7
CodeFeedbackST 59.0 61.2 63.8 74.2 74.9 74.5 74.8 75.5 76.2
CodeFeedbackMT 50.2 49.5 51.7 61.4 71.4 63.6 69.8 70.7 70.4
Mean 43.2 45.9 48.4 52.4 54.5 56.1 55.9 56.4 58.6
Table 11: Retrieval performance (NDCG@10, %) across 13 BEIR tasks for three encoder sizes and three batch sizes (bs4, bs8, bs16).
Revela0.1B Revela0.5B Revela1B
Dataset bs4 bs8 bs16 bs4 bs8 bs16 bs4 bs8 bs16
ArguAna 36.7 38.5 39.0 35.4 41.0 41.1 43.0 47.9 44.6
ClimateFEVER 14.0 15.5 15.3 13.6 15.2 13.8 16.4 14.0 15.8
DBPedia 15.5 15.7 18.3 23.1 20.8 21.3 25.5 20.7 27.6
FEVER 27.0 29.5 33.7 47.3 46.1 51.1 57.7 51.2 61.7
FiQA2018 17.4 18.2 19.2 24.8 27.8 27.2 29.4 28.8 30.4
HotpotQA 32.7 33.3 38.5 45.7 46.0 50.6 51.0 54.6 56.0
NFCorpus 21.1 22.0 23.6 25.0 27.1 26.8 26.6 28.1 27.2
NQ 16.8 17.3 21.2 23.4 27.9 29.8 34.7 32.9 33.9
QuoraRetrieval 68.1 75.8 81.0 83.5 83.6 83.2 83.1 83.0 83.5
SCIDOCS 9.5 10.3 12.1 12.6 14.7 14.8 16.0 15.8 16.3
SciFact 52.4 56.3 57.9 61.1 62.4 66.0 65.5 67.9 71.9
TREC-COVID 58.8 58.1 60.6 62.3 64.1 58.4 67.1 68.8 60.1
Touche2020 14.6 12.7 12.4 21.1 22.4 19.9 26.7 24.7 25.7
Mean 29.6 31.0 33.3 36.8 38.4 38.8 41.7 41.4 42.7
C.3 Analysis on the Size of LMs
Tab. 12 and Tab. 13 present an ablation study analyzing the impact of differently sized LMs.
@@ C.3 Analysis on the Size of LMs
Table 12: NDCG@10 (%) on BEIR for three model sizes (Revela0.1B, Revela0.5B, Revela1B), with LM sizes in ascending order. Each triplet of columns highlights the highest value in bold.
Revela0.1B Revela0.5B Revela1B
@@
ArguAna 36.9 39.9 39.0 38.5 41.8 41.1 41.1 43.0 44.6
C-FEVER 14.1 16.3 15.3 18.3 20.0 13.8 15.0 15.5 15.8
DBPedia 17.6 17.0 18.3 24.2 22.7 21.3 25.0 26.4 27.6
FEVER 25.5 31.3 33.7 53.0 55.7 51.1 58.7 56.7 61.7
FiQA2018 16.9 19.3 19.2 26.6 24.9 27.2 29.3 28.4 30.4
HotpotQA 41.7 41.0 38.5 54.6 53.9 50.6 58.4 58.7 56.0
NFCorpus 19.7 22.2 23.6 27.4 27.4 26.8 26.4 28.7 27.2
NQ 22.3 21.0 21.2 34.5 32.2 29.8 36.8 34.9 33.9
QuoraRetrieval 79.1 79.7 81.0 83.4 83.5 83.2 82.4 83.3 83.5
SCIDOCS 10.8 11.3 12.1 15.4 15.1 14.8 15.9 16.0 16.3
SciFact 52.4 56.7 57.9 65.9 65.8 66.0 71.6 72.1 71.9
TRECCOVID 58.6 60.6 60.6 68.1 68.7 58.4 66.4 65.7 60.1
Touche2020 11.2 13.1 12.4 31.1 24.2 19.9 28.7 26.9 25.7
Mean 31.3 33.0 33.3 41.6 41.2 38.8 42.7 42.8 42.7
@@
Revela0.1B Revela0.5B Revela1B